OpenStudy (anonymous):

Solve for X and Y y=2/3x -3 2x-3y=9

OpenStudy (jwt625):

these two are the same,so we can't get one answer.

OpenStudy (anonymous):

substitue y into the second equation

OpenStudy (anonymous):

but when i substituted them i couldnt find an answer

OpenStudy (anonymous):

x and y can never have values because they are the same witten differently.

OpenStudy (anonymous):

perhaps there is no real solutions i do not know as i have not done the problem only how'd you go about solving a problem like this

OpenStudy (jwt625):

the second equation: 2x-3y=9 2x-9=3y 2/3x-3=y that's the same as the first one.

OpenStudy (jwt625):

so you won't get a concretion answer.

OpenStudy (anonymous):

what Jwt is trying to say is that since they are equal, there will be a infinite amount of solutions